# Modified gravitational wave propagations in linearized gravity with Lorentz and diffeomorphism violations and their gravitational wave constraints

### v1

@@ II.1 Linearized gravity with Lorentz and diffeomorphism violations
where one expands the metric $g_{\mu\nu}$ of the spacetime in the form of $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ with $\eta_{\mu\nu}$ being the constant Minkowski metric, and $\epsilon^{\mu\rho\alpha\kappa}$ is the Levi-Civita tensor. The first term in the above expression represents the quadratic approximation to the Lagrangian density for the Einstein-Hilbert action, while the second term which consists of operators $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ denotes the modifications due to the Lorentz and diffeomorphism violations. The operator $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ is the product of a coefficient ${\cal K}^{(d)\mu\nu\rho\sigma\varepsilon_{1}\varepsilon_{2}\cdots\varepsilon_{d-2}}$ with $d-2$ derivatives $\partial_{\varepsilon_{1}}\partial_{\varepsilon_{2}}\cdots\partial_{\varepsilon_{d-2}}$, i.e.,
$\hat{\cal K}^{(d)\mu\nu\rho\sigma}={\cal K}^{(d)\mu\nu\rho\sigma\varepsilon_{1}\varepsilon_{2}\cdots\varepsilon_{d-2}}\partial_{\varepsilon_{1}}\partial_{\varepsilon_{2}}\cdots\partial_{\varepsilon_{d-2}}.$ (2.2)
The coefficient ${\cal K}^{(d)\mu\nu\rho\sigma\varepsilon_{1}\varepsilon_{2}\cdots\varepsilon_{d-2}}$ has mass dimension $4-d$ and are assumed small and constant over the scales relevant for the gravitational phenomenon considered in this paper. As pointed out in Kostelecký and Mewes (2018), to contribute non-trivially to the equation of motion, the $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ has to satisfy the requirement $\hat{\cal K}^{(d)(\mu\nu)(\rho\sigma)}\neq\pm\hat{\cal K}^{(d)(\rho\sigma)(\mu\nu)}$, where $\pm$ corresponds to odd $d$ and even $d$ respectively.
According to the the symmetries of the coefficient ${\cal K}^{(d)\mu\nu\rho\sigma\varepsilon_{1}\varepsilon_{2}\cdots\varepsilon_{d-2}}$ under the permutations of its indices, $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ can be divided into three different types,
@@
${s}^{(d,0)\mu\rho\alpha_{1}\nu\sigma\alpha_{2}\alpha_{3}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@
${s}^{(d,1)\mu\rho\nu\sigma\alpha_{1}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@
${s}^{(d,2)\mu\rho\alpha_{1}\nu\sigma\alpha_{2}\alpha_{3}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@ II.2 GW propagations with Lorentz- and diffeomorphism-violating effects
In GR, the metric perturbation $h_{\mu\nu}$ contains only two degenerate, traceless, and transverse tensor modes. However, when Lorentz-violating and diffeomorphism-violating modifications are introduced, $h_{\mu\nu}$ may acquire additional degrees of freedom, depending on the specific nature of the violations. These include two scalar modes and two vector modes. As shown in Hou et al. (2024), in Lorentz-violating linearized gravity, the extra scalar and vector modes of gravitational wave (GW) polarization can be directly induced by the two tensorial modes. In this paper, assuming these additional modes are small, and following a similar treatment to that in Kostelecký and Mewes (2016), we focus exclusively on the effects of Lorentz- and diffeomorphism-violating modifications on the two traceless and transverse tensor modes.
@@
Here “$\pm$” correspond to two modes propagating at different velocities. Therefore, the two tensorial modes can be decompose into two modes propagate at different velocities, one is the fast mode (denoted by $h_{f}$ with velocity $v_{+}$) while another is the slow mode (denote by $h_{s}$ with velocity $v_{-}$). One can connect ($h_{f},h_{s}$) with the circular polarization modes ($h_{L},h_{R}$) by normalized matrix as
$\left(\begin{array}[]{ccc}h_{f}\\ h_{s}\end{array}\right)=\left(\begin{array}[]{ccc}-e^{i\varphi/2}\sin\frac{\vartheta}{2}&e^{-i\varphi/2}\cos\frac{\vartheta}{2}\\ e^{i\varphi/2}\cos\frac{\vartheta}{2}&e^{i\varphi/2}\sin\frac{\vartheta}{2}\end{array}\right)\left(\begin{array}[]{ccc}h_{R}\\ h_{L}\end{array}\right).$ (2.38)
Here the angles $\varphi$ and $\vartheta$ are defined as
$\sin\vartheta$ $=$ $\frac{\sqrt{(\zeta^{1})^{2}+(\zeta^{2})^{2}}}{\left|\boldsymbol{\zeta}\right|},$ (2.40)
@@ III Phase Modifications to the Waveform of GWs
Let us turn to derive the modified waveform of GW with Lorentz- and diffeomorphism-violating effects in the linearized gravity sector of the SME. For this purpose, we closely follow the derivation presented in Qiao et al. (2019); Zhao et al. (2020a); Mewes (2019). Now consider a graviton emitted radially at $r=r_{e}$ and received at $r=0$, we have
@@
Let us focus on the GW signal generated by non-spinning, quasi-circular inspiral in the post-Newtonian approximation. Relative to the GW in GR, the Lorentz- and diffeomorphism-violating effects modify the phase of GWs. In Will (1998); Mirshekari et al. (2012), it was proved that the difference of arrival times in (3.3) induces the modification to the phase of GWs $\Psi$ in the following form,
$\Psi(f)=\Psi^{\rm GR}(f)\mp\delta\Psi_{1}(f,{\bf n})+\delta\Psi_{2}(f,{\bf n}),$ (3.4)
@@
Then the fast and slow modes $(h_{f},h_{s})$ with phase corrections can expressed as
@@
$h_{R}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}-i\cos\vartheta\sin\delta\Psi_{1})h_{R}^{GR}$
$~{}~{}~{}~{}~{}~{}~{}~{}-i\sin\vartheta e^{-i\varphi}\sin\delta\Psi_{1}h_{L}^{GR}\Big{]},$ (3.11)
$h_{L}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}+i\cos\vartheta\sin\delta\Psi_{1})h_{L}^{GR}$
$~{}~{}~{}~{}~{}~{}~{}~{}-i\sin\vartheta e^{i\varphi}\sin\delta\Psi_{1}h_{R}^{GR}\Big{]}.$ (3.12)
$h_{+}=\frac{h_{L}+h_{R}}{\sqrt{2}},$ (3.13)
$h_{\times}=\frac{h_{L}-h_{R}}{\sqrt{2i}}.$ (3.14)
@@
$h_{+}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}-i\cos\varphi\sin\vartheta\sin\delta\Psi_{1})h_{+}^{GR}$
$~{}~{}~{}~{}~{}~{}~{}-(\cos\vartheta+i\sin\vartheta\sin\varphi)\sin\delta\Psi_{1}h_{\times}^{GR}\Big{]},$ (3.15)
$h_{\times}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}+i\cos\varphi\sin\vartheta\sin\delta\Psi_{1})h_{\times}^{GR}$
$~{}~{}~{}~{}~{}~{}~{}+(\cos\vartheta-i\sin\vartheta\sin\varphi)\sin\delta\Psi_{1}h_{+}^{GR}\Big{]}.$ (3.16)
Here we would like to consider several special limits of the above general waveform. Considering that the operators with the lowest mass dimension are expected to have the dominant Lorentz- and diffeomorphism-violating effects on the propagation of GWs, in the following we only discuss specific cases with relative lower mass dimensions, for example, $d=2,3,4,5,6$. Note that ref. Gong et al. (2022) explores isotropic Lorentz-violating effects on GWs, which correspond to operators with mass dimensions $d=7$ and $d=8$.

### v2

@@ II.1 Linearized gravity with Lorentz and diffeomorphism violations
where one expands the metric $g_{\mu\nu}$ of the spacetime in the form of $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ with $\eta_{\mu\nu}$ being the constant Minkowski metric, and $\epsilon^{\mu\rho\alpha\kappa}$ is the Levi-Civita tensor. The first term in the above expression represents the quadratic approximation to the Lagrangian density for the Einstein-Hilbert action, while the second term which consists of operators $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ denotes the modifications due to the Lorentz and diffeomorphism violations. The operator $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ is the product of a coefficient ${\cal K}^{(d)\mu\nu\rho\sigma{\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}}$ with $d-2$ derivatives $\partial_{\alpha_{1}}\partial_{\alpha_{2}}\cdots\partial_{\alpha_{d-2}}$, i.e.,
$\hat{\cal K}^{(d)\mu\nu\rho\sigma}={\cal K}^{(d)\mu\nu\rho\sigma\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}{\partial_{\alpha_{1}}\partial_{\alpha_{2}}\cdots\partial_{\alpha_{d-2}}}.$ (2.2)
The coefficient ${\cal K}^{(d)\mu\nu\rho\sigma\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}$ has mass dimension $4-d$ and are assumed small and constant over the scales relevant for the gravitational phenomenon considered in this paper. As pointed out in Kostelecký and Mewes (2018), to contribute non-trivially to the equation of motion, the $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ has to satisfy the requirement $\hat{\cal K}^{(d)(\mu\nu)(\rho\sigma)}\neq\pm\hat{\cal K}^{(d)(\rho\sigma)(\mu\nu)}$, where $\pm$ corresponds to odd $d$ and even $d$ respectively.
According to the the symmetries of the coefficient ${\cal K}^{(d)\mu\nu\rho\sigma{\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}}$ under the permutations of its indices, $\hat{\cal K}^{(d)\mu\nu\rho\sigma}$ can be divided into three different types,
@@
$\hat{s}^{(d,0)\mu\rho\alpha_{1}\nu\sigma\alpha_{2}\alpha_{3}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@
$\hat{s}^{(d,1)\mu\rho\nu\sigma\alpha_{1}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@
$\hat{s}^{(d,2)\mu\rho\alpha_{1}\nu\sigma\alpha_{2}\alpha_{3}\ldots\alpha_{d-2}}$ytableausetupcentertableaux,boxsize=2.5em
@@ II.2 GW propagations with Lorentz- and diffeomorphism-violating effects
In GR, the metric perturbation $h_{\mu\nu}$ contains only two degenerate, traceless, and transverse tensor modes. However, when Lorentz-violating and diffeomorphism-violating modifications are introduced, $h_{\mu\nu}$ may acquire additional degrees of freedom, depending on the specific nature of the violations. These include two scalar modes and two vector modes. As shown in Hou et al. (2024), in Lorentz-violating linearized gravity, the extra scalar and vector modes of gravitational wave (GW) polarization can be directly induced by the two tensorial modes. Here in this paper we restrict our study to the two transverse and traceless modes of GWs. Observationally, all the GW signals observed by the LIGO/Virgo/KAGRA collaboration exhibit compatibility with the two tensor polarization modes, showing no statistically significant signatures of additional polarization modes Abbott et al. (2023). Even if additional modes exist, their effects are expected to be be subdominant compared to the two tensorial modes. In this paper, assuming these additional modes are small, and following a similar treatment to that in Kostelecký and Mewes (2016), we focus exclusively on the effects of Lorentz- and diffeomorphism-violating modifications on the two traceless and transverse tensor modes. We note that this assumption imposes certain limitations, as our conclusions may not fully account for scenarios where such subdominant modes play a more substantial role. We left the analysis with the vector and scalar modes for future studies.
@@
Here “$\pm$” correspond to two modes propagating at different velocities. Therefore, the two tensorial modes can be decompose into two modes propagate at different velocities, one is the fast mode (denoted by $h_{f}$ with velocity $v_{+}$) while another is the slow mode (denote by $h_{s}$ with velocity $v_{-}$).
Now we need to connect ($h_{f},h_{s}$) with the circular polarization modes ($h_{L},h_{R}$). For this purpose, one can substitute the dispersion relation in Eq. (2.25) to the equation of motion (2.24) for fast and slow modes respectively. Note that in the substitution, we use the relation $\partial_{t}^{2}=-\omega^{2}$. For fast mode, by using the experssions of $\zeta_{0}$ and ${\boldsymbol{\zeta}}$ in Eqs. (II.2) and  (2.27), one has
$0\displaystyle 2|{\bf k}|^{2}\left(\begin{array}[]{cc}\zeta^{3}-\left|\boldsymbol{\zeta}\right|&\zeta^{1}-i\zeta^{2}\\ \zeta^{1}+i\zeta^{2}&-\zeta^{3}-\left|\boldsymbol{\zeta}\right|\end{array}\right)\left(\begin{array}[]{c}h_{R}\\ h_{L}\end{array}\right)=$ 0 . (2.36)
This equation admits a set of solution in the form of
$\left(\begin{array}[]{c}h_{R}\\ h_{L}\end{array}\right)_{\rm fast}=\left(\begin{array}[]{c}e^{-i\varphi/2}\cos\frac{\vartheta}{2}\\ e^{i\varphi/2}\sin\frac{\vartheta}{2}\end{array}\right),$ (2.41)
where the angles $\varphi$ and $\vartheta$ are defined as
$\sin\vartheta$ $=$ $\frac{\sqrt{(\zeta^{1})^{2}+(\zeta^{2})^{2}}}{\left|\bm{\zeta}\right|},$ (2.42)
@@
Similarly, for slow mode, one has
$\left(\begin{array}[]{c}h_{R}\\ h_{L}\end{array}\right)_{\rm slow}=\left(\begin{array}[]{c}-e^{-i\varphi/2}\sin\frac{\vartheta}{2}\\ e^{i\varphi/2}\cos\frac{\vartheta}{2}\end{array}\right).$ (2.49)
Therefore, Eq. (2.41) and (2.49) can be written as a unique equation as
${\left(\begin{array}[]{ccc}h_{R}\\ h_{L}\end{array}\right)=\left(\begin{array}[]{ccc}e^{-i\varphi/2}\cos\frac{\vartheta}{2}&-e^{-i\varphi/2}\sin\frac{\vartheta}{2}\\ e^{i\varphi/2}\sin\frac{\vartheta}{2}&e^{i\varphi/2}\cos\frac{\vartheta}{2}\end{array}\right)\left(\begin{array}[]{ccc}h_{f}\\ h_{s}\end{array}\right).}$ (2.56)
It is also convenient to express $(h_{f},h_{s})$ in terms of $(h_{R},h_{L})$ as
$\left(\begin{array}[]{ccc}h_{f}\\ h_{s}\end{array}\right)=\left(\begin{array}[]{ccc}e^{i\varphi/2}\cos\frac{\vartheta}{2}&e^{-i\varphi/2}\sin\frac{\vartheta}{2}\\ -e^{i\varphi/2}\sin\frac{\vartheta}{2}&e^{-i\varphi/2}\cos\frac{\vartheta}{2}\end{array}\right)\left(\begin{array}[]{ccc}h_{R}\\ h_{L}\end{array}\right).$ (2.64)
$h_{+}=\frac{h_{L}+h_{R}}{\sqrt{2}},$ (2.66)
$h_{\times}=\frac{h_{L}-h_{R}}{\sqrt{2}i}.$ (2.67)
Then we can write $(h_{f},h_{s})$ in terms of $h_{+}$ and $h_{\times}$ as
$\begin{pmatrix}h_{f}\\ h_{s}\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}-e^{i\varphi/2}\sin\frac{\vartheta}{2}+e^{-i\varphi/2}\cos\frac{\vartheta}{2}&ie^{i\varphi/2}\sin\frac{\vartheta}{2}+ie^{-i\varphi/2}\cos\frac{\vartheta}{2}\\ e^{i\varphi/2}\cos\frac{\vartheta}{2}+e^{-i\varphi/2}\sin\frac{\vartheta}{2}&-ie^{i\varphi/2}\cos\frac{\vartheta}{2}+ie^{-i\varphi/2}\sin\frac{\vartheta}{2})\end{pmatrix}\begin{pmatrix}h_{+}\\ h_{\times}\end{pmatrix}.$ (2.68)
And inversely,
${\begin{pmatrix}h_{+}\\ h_{\times}\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}e^{i\varphi/2}\cos\frac{\vartheta}{2}-e^{-i\varphi/2}\sin\frac{\vartheta}{2}&e^{-i\varphi/2}\cos\frac{\vartheta}{2}+e^{i\varphi/2}\sin\frac{\vartheta}{2}\\ -ie^{-i\varphi/2}\sin\frac{\vartheta}{2}-ie^{i\varphi/2}\cos\frac{\vartheta}{2}&-ie^{i\varphi/2}\sin\frac{\vartheta}{2}+ie^{-i\varphi/2}\cos\frac{\vartheta}{2}\end{pmatrix}\begin{pmatrix}h_{f}\\ h_{s}\end{pmatrix}.}$ (2.69)
@@ III Phase Modifications to the Waveform of GWs
In this section, we consider the propagation of GWs with Lorentz- and diffeomorphism-violating effects in a Friedman-Robertson-Walker (FRW) background, whose metric is given by,
$ds^{2}=-dt^{2}+a^{2}(t)d\mathbf{r}^{2},$ (3.1)
where $t$ is the cosmic time, $a(t)$ is the scale factor governing the expansion of the universe. Now we consider a graviton emitted radially at $r=r_{e}$ and received at $r=0$, we have
@@
Let us focus on the GW signal generated by non-spinning, quasi-circular inspiral in the post-Newtonian approximation. Relative to the GW in GR, the Lorentz- and diffeomorphism-violating effects modify the phase of GWs. The phase corrections to the GR-based waveform due to Lorentz- and diffeomorphism-violating effects can be computed using the stationary phase approximation (SPA) during the inspiral phase of the binary system Zhao et al. (2020a); Mirshekari et al. (2012). As demonstrated in Ezquiaga et al. (2022), the waveforms modified by propagation effects and derived using the SPA agree with those derived using the WKB approximation. In WKB approximation, the corrections to the GR-based waveform are only due to the propagation effect, and thus it is in principle independent of the GW emission mechanism or radiated stages of the binary system Ezquiaga et al. (2022). This implies that one can extend the modified waveforms obtained using the SPA Zhao et al. (2020a); Mirshekari et al. (2012) to the entire signal including the inspiral, merger, and ringdown phases of a coalescing binary system.
In the SPA, the Fourier transform of $h_{A}(t)$ can be obtained analytically, which is given by Maggiore (2007)
$\tilde{h}_{A}(f)=\frac{{\cal A}_{A}(f)}{\sqrt{df/dt}}e^{i\Psi(f)},$ (3.5)
where $f$ is the GW frequency at the detector, ${\cal A}_{A}(f)$ is the amplitude, and $\Psi$ is the phase of GWs. In Will (1998); Mirshekari et al. (2012), it was proved that the difference of arrival times in (3.4) induces the modification to the phase of GWs $\Psi$ in the following form,
$\Psi(f)=\Psi^{\rm GR}(f)+\delta\Psi(f),$ (3.6)
where $\Psi^{\rm GR}(f)$ is the phase predicted in GR and $\delta\Psi(f)$ is the phase corrections due to the Lorentz- and diffeomorphism-violating effects. According to the time difference in Eq. (3.4), the phase corrections can be divided into two parts,
$\delta\Psi(f)=\mp\delta\Psi_{1}(f,{\bf n})+\delta\Psi_{2}(f,{\bf n}),$ (3.7)
where $\delta\Psi_{1}(f,{\bf n})$ corresponds to the velocity birefringence effect and $\delta\Psi_{2}(f,{\bf n})$ represents the non-birefringent effects. $\delta\Psi_{1}$ is induced by the Lorentz- and diffeomorphism-violating coefficients $k^{(d)}_{(V)jm}$ and $k^{(E)}_{(I)jm}\pm ik^{(d)}_{(B)jm}$, while $\delta\Psi_{2}$ is induced by $k^{(d)}_{(I)jm}$.
@@
In this paper, we adopt the above modified waveforms for later analysis with the open data of compact binary merging events detected by the LIGO-Virgo-KAGRA collaboration.
Let us turn to derive the modified waveform of GW with Lorentz- and diffeomorphism-violating effects in the linearized gravity sector of the SME. For this purpose, we closely follow the derivation presented in Qiao et al. (2019); Zhao et al. (2020a); Mewes (2019).Then the fast and slow modes $(h_{f},h_{s})$ with phase corrections can expressed as
@@
$h_{R}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}-i\cos\vartheta\sin\delta\Psi_{1})h_{R}^{\text{GR}}$ (3.15)
$~{}~{}~{}~{}~{}~{}~{}~{}-i\sin\vartheta e^{-i\varphi}\sin\delta\Psi_{1}h_{L}^{\text{GR}}\Big{]},$
$h_{L}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}+i\cos\vartheta\sin\delta\Psi_{1})h_{L}^{\text{GR}}$
$~{}~{}~{}~{}~{}~{}~{}~{}-i\sin\vartheta e^{i\varphi}\sin\delta\Psi_{1}h_{R}^{\text{GR}}\Big{]}.$
@@
$h_{+}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}-i\cos\varphi\sin\vartheta\sin\delta\Psi_{1})h_{+}^{\text{GR}}$
$~{}~{}~{}~{}~{}~{}~{}-(\cos\vartheta+i\sin\vartheta\sin\varphi)\sin\delta\Psi_{1}h_{\times}^{\text{GR}}\Big{]},$ (3.16)
$h_{\times}=e^{i\delta\Psi_{2}}\Big{[}(\cos\delta\Psi_{1}+i\cos\varphi\sin\vartheta\sin\delta\Psi_{1})h_{\times}^{\text{GR}}$
$~{}~{}~{}~{}~{}~{}~{}+(\cos\vartheta-i\sin\vartheta\sin\varphi)\sin\delta\Psi_{1}h_{+}^{\text{GR}}\Big{]}.$ (3.17)
Here we would like to add two remarks about the above modified waveforms and their tests with GW signals detected by LIGO-Virgo-KAGRA detectors. Firstly, the modified waveform presented in the above represents the most general waveform with Lorentz- and diffeomorphism-violating effects in the linearized gravity. Here we note that the modified waveforms with Lorentz-violating but diffeomorphism invariant coefficients for $d>4$ has been considered in refs. Gong et al. (2023); Kostelecký and Mewes (2016); Haegel et al. (2023); Niu et al. (2022); O’Neal-Ault et al. (2021). The waveforms with phase corrections presented above generalize those in Refs. Gong et al. (2023); Kostelecký and Mewes (2016); Haegel et al. (2023); Niu et al. (2022); O’Neal-Ault et al. (2021) to the Lorentz- and diffeomorphism-violating case, valid for $d\geq 2$. Secondly, the constraints on the Lorentz-violating coefficients with mass dimensions $d=5$ and $d=6$ have been studied by comparing the above modified waveforms with GW signals detected by LIGO-Virgo-KAGRA detectors have been performed in refs. Gong et al. (2023); Haegel et al. (2023); Niu et al. (2022). And the main purpose of the rest part of this paper is to use the above modified waveforms to derive the constraints on the Lorentz- and diffeomorphism-violating coefficients for $d=2$ and $d=3$ cases.
Then we would like to consider several special limits of the above general waveform. Considering that the operators with the lowest mass dimension are expected to have the dominant Lorentz- and diffeomorphism-violating effects on the propagation of GWs, in the following we only discuss specific cases with relative lower mass dimensions, for example, $d=2,3,4,5,6$. Note that ref. Gong et al. (2022) explores isotropic Lorentz-violating effects on GWs, which correspond to operators with mass dimensions $d=7$ and $d=8$.